40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 1966     Entry time: Thu Sep 3 23:41:32 2009
Author: Alberto 
Type: Configuration 
Category: LSC 
Subject: POX (PD3) aligned 

Today I aligned the beam to PD3 (POX) since Steve had moved it.

The DC power read 1.3mV when the beam was on the PD.

ELOG V3.1.3-